OpenStudy (anonymous):

HELP ME? Show me how to do the work also - Calculate the speed of an 8.0 x 10^4 kg airliner with a kinetic energy of 1.1 x 10^9 J.

OpenStudy (anonymous):

Above my pay grade, but don't you need a drag coefficient and altitude/atmospheric pressure to calculate this? They both effect speed.

OpenStudy (anonymous):

Or is this equation, "static in a vacuum"

OpenStudy (anonymous):

we are assuming no external forces. let K = kinetic energy, m = mass, V = velocity/speed \[K = \frac{ 1 }{ 2 }m V^2\] \[1.1 * 10^9 = \frac{ 1 }{ 2 }8.0*10^4* V^2\] solve for V

OpenStudy (anonymous):

I got this \[1.1 x 10^9 = 1/2*8.0*10^4\] \[1100000000=0.5*8000\] \[1100000000 = 4000\] \[524.4 \] Did I do this right? Answers say its 1.7 x 10^2? Am I right?

OpenStudy (anonymous):

i get 1.658* 10^2

OpenStudy (anonymous):

\[V = \sqrt{\frac{ 1,100,000,000 }{ 0.5*80,000 }} = \sqrt{27500}\] but you should really enter 1.1 E 9 in your calculator (1.1 * 10^9) if you can't find the button, i can help.

OpenStudy (anonymous):

Thank you, I understand now(: Could you help me with another problem?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

What is the speed of a 0.145 kg baseball if the kinetic energy is 109 J? I used the same equation and got 31.61 . I think I made a mistake. Answers say 38.8

OpenStudy (anonymous):

i get 38.77 \[K = \frac{ 1 }{ 2 }m V^2\] \[V = \sqrt{\frac{ 2K }{ m }} = \sqrt{\frac{ 2*109 }{ 0.145 }}\]

OpenStudy (anonymous):

How did you get the 2K? Thats the part I don't understand?

OpenStudy (anonymous):

i'm just isolating V from:\[K = \frac{ 1 }{ 2 }m V^2\] multiply both sides by 2:\[2K = mV^2\]divide both sides by m\[\frac{ 2K }{ m } = V^2\]

OpenStudy (anonymous):

what level calc is this problem from? wondering when I can look forward to these equations.

OpenStudy (anonymous):

he multiplied both sides by 2 to clear the fraction

OpenStudy (anonymous):

this equation is from mechanics or general physics. you will most likely be taking this class in the same time as cal 1

OpenStudy (anonymous):

I can't take physics or mech until concurrent with calc 2

OpenStudy (anonymous):

that is different that here ^_^. probably means you will be going straight into applying integration to basic physics. early physics classes here don't require calculus

OpenStudy (anonymous):

I need help with two more problems if thats okay?